# Probabilistic-Numeric SMC Sampling for Bayesian Nonlinear System Identification in Continuous Time

### v1

@@ 1 Introduction
where, $f$ and $g$ govern the transitions of the state ($x$) and observation $y$ respectively, parameterised by $\theta_{f}$ and $\theta_{g}$. In the parameter identification setting, it is assumed that the functional form of $f$ and $g$ are known.
@@
where $x_{t}\triangleq x(t)$ and $p(x_{0})$ denotes the prior probability distribution of the initial state, encapsulating a practitioners knowledge before any measurements are taken. The transition distribution $p(x_{t}\mid x_{t-1})$ characterises the system dynamics and inherent uncertainties, describing the evolution of the state from $x_{t-1}$ to $x_{t}$. Lastly, the measurement model $p(y_{t}\mid x_{t})$ represents the relationship between the current state $x_{t}$ and the measurement $y_{t}$, reflecting how the observed data arises from the underlying state of the system.
@@ 1.1 Contribution
The contribution of this work is to a nonlinear system identification method which combines the ideas of PN for solving ODEs — as (nonlinear) filtering problems where the uncertainty in the numerical integration is quantified — with an SMC scheme which iteratively approaches the posterior over the model parameters. The strength of this approach is that uncertainty from both measurement noise and the numerical methods employed is respected in the inference over the model paramters. Additionally, since SMC is a recursive algorithm, each intermediate posterior is a valid posterior over the parameters conditioned on an increasing subset of the time-series data; consequently, this approach provides a route to online Bayesian estimation of nonlinear system parameters (subject to the computational requirements at each step being achieveable within one sampling period). Finally, by explicitly including the numerical integration of the ODEs within the parameter identification procedure the methodolgy presented seeks to directly identify the continuous time ODE(s) which govern the nonlinear system.
@@ 2.1 Background
The simulation and analysis of dynamic systems are often rooted in the concept of Initial Value Problems (IVPs), typically presented as solving Ordinary Differential Equations (ODEs) with known initial conditions:
$\dot{{v}}=g(t,v(t)),\quad v(t_{0})=v_{0},$ (5)
where $v(t)\in\mathbb{R}^{d}$ is the solution, $g(\cdot)$ is a function that maps the current states to their respective derivatives, and $v_{0}$ denotes the initial states at time $t_{0}$. In a broader context, engineers are often interested in dynamic systems under forced excitation. The SSM transition function $f(\cdot)$ can be used to extend the IVP framework to accommodate forced excitation:
@@ 2.2 From Ordinary to Stochastic Differential Equations
where ${\mu}(1)$ and $\Sigma(1)$ are the mean and the covariance that describe the Gaussian distribution over the initial conditions ${X}(1)$. $F$ is the state transition matrix, $u(t)$ the force, $L$ is the diffusion matrix and $\beta(t)$ is defined as a vector of the standard Wiener processes. ${X}(t)$ is a vector of $X^{(1)}(t)$ and $q+1$ derivatives such that $X^{q+1}(t)$ is the derivative of $X^{q}(t)$,
@@
To fit the SDE in a Bayesian filtering framework measurements are required. For solving IVPs, Tronarp [17] proposes that a measurement can be defined upon the known derivative relationship between the states i.e. $F$, $u$ and $L$ must be set such that $X^{q+1}(t)$ is always be the derivative of $X^{q}(t)$. This provides the following pseudo-measurement
@@ 2.3 The Discrete Time Solution
where $h$ is the step size and $R$ represents the measurement variance. In this work and in general when solving Eq.(12) with Gaussian filtering $R=0$. $C$ is the observation matrix and $\dot{C}$ is it’s derivative such that $C=\begin{bmatrix}\mathbf{I}&0&\cdots&0\end{bmatrix}$ and $\dot{C}=\begin{bmatrix}0&\mathbf{I}&\cdots&0\end{bmatrix}$. That is, ${C}{X}_{t}={X}_{t}^{(1)}$ and $\dot{C}{X}_{t}={X}_{t}^{(2)}$. It is important to observe that the likelihood model $p(Z_{t}\mid X_{t})$ exhibits nonlinearity, which adds complexity to the filtering solution. Additionally, $z_{t}$ denotes the realization of $Z_{t}$, and the state transition $A(h)$, process noise $\xi(h)$, and process noise covariance $Q(h)$ are defined as follows:
@@ 2.6 Calibration
However, computing this is as costly as solving the ODE, thus necessitating an approximation. Bosch et al. [26] proposed a quasi-ML estimator for multidimensional ODEs in tha case of the EKF0, expressed as:
@@
where $S_{t}=\Gamma\breve{s}_{t}$ and $S_{t}$ is the innovation covariance matrix from the EKF0 equations 17. This estimator, while an approximation, offers efficient calibration for the variance of the posterior and can be easily Incorporated into the EKF0 for almost no additional cost.
@@ 3.1 Sequential Monte Carlo for parameter estimation
9:        Resample a new particle system $\{\tilde{\theta}_{m},\tilde{w}_{m}\}_{m=1:N}$ using an Independent Metropolis Hastings proposal, Eq.(23)
@@
The algorithm initiates at $t_{1}$, where the parameter distribution is represented by the prior $p(\theta)$, as informed by the practitioner’s engineering knowledge. Direct conditioning of $p(\theta)$ on $y_{1:T}$ is computationally intractable. Therefore, $p(\theta|y_{1:T})$ will be approximated by conditioning on a discrete set of random samples $\theta_{1:N}\sim p(\theta)$ (Algorithm 1.1). Using these sampled particles, initial conditions $X_{t_{1},1:N}$, and the probabilistic numerical integrator (referenced in Section 2), the distribution over the system states at subsequent time steps $({X}_{t+1,n}\mid{Z}_{t},{\theta}_{n})_{n=1:N}$ can be determined (Algorithm 1.3).
@@
For the move kernel to have a good efficiency it is important that the acceptance rate of new parameters it high. This could be easily achieved artificially via a random walk where a small perturbation to current particles is applied. However, this fails to actually rejuvenate the particles as although particles are distinct they are also very similar and therefore very highly correlated. This approach leaves degeneracy high but less detectable.
Therefore, a move kernel should be chosen that only proposes new particles that weakly depend on the previous values. For this independent Metropolis Hasting (IMH) kernel is chosen [10]. This ensures that the acceptance rate becomes better indicator of rejuvenation.
Since the $p(\theta|y_{1:t},Z_{1:t})$ is approximated by the weighted particles, resampling and rejuvenation can be performed by sampling from that distribution. A rough approximation of the location of the mass of $p(\theta\mid y_{1:t},Z_{1:t}))$ is given by the expectation ($\hat{E}$) and variance ($\hat{V}$) of the particle system,
@@
The employment of the IMH kernel within SMC methods serves a dual purpose: resampling and rejuvenation of the particle set. This dual functionality stems from the kernel’s capability to sample from the posterior distribution $p(\theta|y_{1:t}Z_{1:t})$ as approximated by the weighted particles. Such sampling not only ensures that the distribution of particles closely approximates the posterior distribution $p(\theta|y_{1:t}Z_{1:t}))$ itself but also facilitates an effective exploration of the parameter space by the particles. Finally, the newly resampled and rejuvenated particle system is input back into the outer loop and the particle weights are reset to $1/N$ so that the process of evaluation given new observations can continue until rejuvenation is once again required. This process is repeated for each of the the discrete time steps in the data set (Algorithm 1.12).
@@ 4.1 Bouc Wen
Figure 1: Prior and posterior hystograms and PDFs normalised by the true parameter values for the Bouc Wen system (denoted by $(\cdot)^{*}$) .
@@ 4.2 Silverbox
Figure 4: Measured input and resonce of the silver box benchmark.
@@
Figure 5: Prior and posterior hystograms and PDFs for the silver box system.
@@
Since, no ground truth is available to evaluate the quality of the identified parameters an alternative method must be used. For this the identified parameters will be used in simulation and compared to the observations for the for the arrowhead section of the dataset. The first 1000 time steps were excluded from this simulation to remove the transient so that the simulation was run from time step 1000 to 40,000. For this simulation a RMSE was calculated for each particle. The maximum RMSE was 2.9516$\times$10-3 the minimum was 1.0567$\times$10-3 and the mean RMSE of all particles was 1.8249$\times$10-3.
@@ 5 Conclusions
It is the argument of the authers that: parameter estimation in nonlinear dynamic systems can abstractly be delineated into 3 sequential stages. Step 1 is to select a set of trial parameters. Step 2 involves using the set of trial parameters, the functional form of the model, and initial conditions to solve the IVP. Solving the IVP gives access to the state of the system at some future point in time. Step three is to compare the state pertaining to the IVP to a measured state at the same instance in time. By minimising the distance between the proposed and measured state it is possible to optimise the model parameters.
@@ Acknowledgements
The authors gratefully acknowledge the support of the Engineering and Physical Sciences Research Council (EPSRC), UK through grant number EP/W002140/1 and EP/L016257/1. The authors would also like to thank Ramboll Energy for their ongoing support. For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.

### v2

# Probabilistic Numeric SMC Sampling for Bayesian Nonlinear System Identification in Continuous Time

@@ 1 Introduction
where, $f$ and $g$ govern the transitions of the state $x$ and observation $y$ respectively, parameterised by $\theta_{f}$ and $\theta_{g}$. In the parameter identification setting, it is assumed that the functional form of $f$ and $g$ are known.
@@
where $x_{t}\triangleq x(t)$ and $p(x_{1})$ denotes the prior probability distribution of the initial state, encapsulating a practitioners knowledge before any measurements are taken. The transition distribution $p(x_{t}\mid x_{t-1})$ characterises the system dynamics and inherent uncertainties, describing the evolution of the state from $x_{t-1}$ to $x_{t}$. Lastly, the measurement model $p(y_{t}\mid x_{t})$ represents the relationship between the current state $x_{t}$ and the measurement $y_{t}$, reflecting how the observed data arises from the underlying state of the system.
@@ 1.1 Contribution
The contribution of this work is to a nonlinear system identification method which combines the ideas of PN for solving ODEs — as (nonlinear) filtering problems where the uncertainty in the numerical integration is quantified — with an SMC scheme which iteratively approaches the posterior over the model parameters. The strength of this approach is that uncertainty from both measurement noise and the numerical methods employed is respected in the inference over the model parameters. Additionally, since SMC is a recursive algorithm, each intermediate posterior is a valid posterior over the parameters conditioned on an increasing subset of the time-series data; consequently, this approach provides a route to online Bayesian estimation of nonlinear system parameters (subject to the computational requirements at each step being achievable within one sampling period). Finally, by explicitly including the numerical integration of the ODEs within the parameter identification procedure the methodology presented seeks to directly identify the continuous time ODE(s) which govern the nonlinear system.
@@ 2.1 Background
The simulation and analysis of dynamic systems are often rooted in the concept of Initial Value Problems (IVPs), typically presented as solving ODEs with known initial conditions:
$\dot{{v}}=g(t,v(t)),\quad v(t_{1})=v_{1},$ (5)
where $v(t)\in\mathbb{R}^{d}$ is the solution, $g(\cdot)$ is a function that maps the current states to their respective derivatives, and $v_{1}$ denotes the initial states at time $t_{1}$. In a broader context, engineers are often interested in dynamic systems under forced excitation. The SSM transition function $f(\cdot)$ can be used to extend the IVP framework to accommodate forced excitation:
@@ 2.2 From Ordinary to Stochastic Differential Equations
where ${\mu}(1)$ and $\Sigma(1)$ are the mean and the covariance that describe the Gaussian distribution over the initial conditions ${X}(1)$. $F$ is the state transition matrix, $u(t)$ the force, $L$ is the diffusion matrix and $\beta(t)$ is defined as a vector of the standard Wiener processes. ${X}(t)$ is a vector of $X^{(1)}(t)$ and $q+1$ derivatives such that $X^{i+1}(t)$ is the derivative of $X^{i}(t)$,
@@
To fit the SDE in a Bayesian filtering framework measurements are required. For solving IVPs, Tronarp [17] proposes that a measurement can be defined upon the known derivative relationship between the states i.e. $F$, $u$ and $L$ must be set such that $X^{i+1}(t)$ is always be the derivative of $X^{i}(t)$. This provides the following pseudo-measurement
@@ 2.3 The Discrete Time Solution
where $h$ is the step size and $R$ represents the measurement variance. In this work and in general when solving Eq.(12) with Gaussian filtering $R=0$. $C$ is the observation matrix and $\dot{C}$ is it’s derivative such that $C=\begin{bmatrix}\mathbf{I}&0&\cdots&0\end{bmatrix}$ and $\dot{C}=\begin{bmatrix}0&\mathbf{I}&\cdots&0\end{bmatrix}$. That is, ${C}{X}_{t}={X}_{t}^{(1)}$ and $\dot{C}{X}_{t}={X}_{t}^{(2)}$. It is important to observe that the likelihood model $p(Z_{t}\mid X_{t})$ exhibits nonlinearity, which adds complexity to the filtering solution. Additionally, $z_{t}$ denotes the realisation of $Z_{t}$, and the state transition $A(h)$, process noise $\xi(h)$, and process noise covariance $Q(h)$ are defined as follows:
@@ 2.6 Calibration
However, computing this is as costly as solving the ODE, thus necessitating an approximation. Bosch et al. [26] proposed a quasi-ML estimator for multidimensional ODEs in the case of the EKF0, expressed as:
@@
where $S_{t}=\Gamma\breve{s}_{t}$ and $S_{t}$ is the innovation covariance matrix from the EKF0 equations Eq.(17). This estimator, while an approximation, offers efficient calibration for the variance of the posterior and can be easily Incorporated into the EKF0 for almost no additional cost.
@@ 3.1 Sequential Monte Carlo for parameter estimation
9:        Sample a new particle system $\{\tilde{\theta}_{m},\tilde{w}_{m}\}_{m=1:N}$ using an Independent Metropolis Hastings proposal, Eq.(23)
@@
The algorithm initiates at $t_{1}$, where the parameter distribution is represented by the prior $p(\theta)$, as informed by the practitioner’s engineering knowledge. Direct conditioning of $p(\theta)$ on $y_{1:T}$ is computationally intractable. Therefore, $p(\theta|y_{1:T})$ will be approximated by conditioning on a discrete set of random samples $\theta_{1:N}\sim p(\theta)$ (Algorithm 1.1). Using these sampled particles, initial conditions $X_{t_{1},1:N}$, and the probabilistic numerical integrator (referenced in Section 2), the distribution over the system states at subsequent time steps $({X}_{t+1}\mid{Z}_{t},{\theta}_{n})_{n=1:N}$ can be determined (Algorithm 1.3).
@@
For the move kernel to have a good efficiency it is important that the acceptance rate of new parameters is high. This could be easily achieved artificially via a random walk where a small perturbation to current particles is applied. However, this fails to actually rejuvenate the particles as although particles are distinct they are also very similar and therefore very highly correlated. This approach leaves degeneracy high but less detectable.
Therefore, a move kernel should be chosen that only proposes new particles that weakly depend on the previous values. For this the independent Metropolis Hasting (IMH) kernel is chosen [10]. This ensures that the acceptance rate becomes better indicator of rejuvenation.
Since the $p(\theta|y_{1:t},Z_{1:t})$ is approximated by the weighted particles, resampling and rejuvenation can be performed by sampling from that distribution. A rough approximation of the location of the mass of $p(\theta\mid y_{1:t},Z_{1:t})$ is given by the expectation $\hat{E}$ and variance $\hat{V}$ of the particle system,
@@
The employment of the IMH kernel within SMC methods serves a dual purpose: resampling and rejuvenation of the particle set. This dual functionality stems from the kernel’s capability to sample from the posterior distribution $p(\theta|y_{1:t}Z_{1:t})$ as approximated by the weighted particles. Such sampling not only ensures that the distribution of particles closely approximates the posterior distribution $p(\theta|y_{1:t}Z_{1:t})$ itself but also facilitates an effective exploration of the parameter space by the particles. Finally, the newly resampled and rejuvenated particle system is input back into the outer loop and the particle weights are reset to $1/N$ so that the process of evaluation given new observations can continue until rejuvenation is once again required. This process is repeated for each of the the discrete time steps in the data set (Algorithm 1.12).
@@ 4.1 Bouc Wen
Figure 1: Prior and posterior histograms and PDFs normalised by the true parameter values for the Bouc Wen system (denoted by $(\cdot)^{*}$) .
@@ 4.2 Silverbox
Figure 4: Measured input and response of the silver box benchmark.
@@
Figure 5: Prior and posterior histograms and PDFs for the silver box system.
@@
Since, no ground truth is available to evaluate the quality of the identified parameters an alternative method must be used. For this the identified parameters will be used in simulation and compared to the observations for the arrowhead section of the dataset. The first 1000 time steps were excluded from this simulation to remove the transient so that the simulation was run from time step 1000 to 40,000. For this simulation a RMSE was calculated for each particle. The maximum RMSE was 2.9516$\times$10-3 the minimum was 1.0567$\times$10-3 and the mean RMSE of all particles was 1.8249$\times$10-3.
@@ 5 Conclusions
It is the argument of the authors that: parameter estimation in nonlinear dynamic systems can abstractly be delineated into 3 sequential stages. Step 1 is to select a set of trial parameters. Step 2 involves using the set of trial parameters, the functional form of the model, and initial conditions to solve the IVP. Solving the IVP gives access to the state of the system at some future point in time. Step 3 is to compare the state pertaining to the IVP to a measured state at the same instance in time. By minimising the distance between the proposed and measured state it is possible to optimise the model parameters.
@@ Acknowledgements
The authors gratefully acknowledge the support of the Engineering and Physical Sciences Research Council (EPSRC), UK through grant number EP/W002140/1 and EP/W005816/1. For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.